OpenStudy (anonymous):

This exercise uses the radioactive decay model. A wooden artifact from an ancient tomb contains 80% of the carbon-14 that is present in living trees. How long ago was the artifact made? (The half-life of carbon-14 is 5730 years. Round your answer to the nearest whole number.) I need the answer in yrs.

OpenStudy (anonymous):

This is what I come up with: \[N(t)=N _{0}(.5)^{t/t _{half-life}} Insert values \rightarrow .8N _{0}=N _{0}(.5)^{t/5730}\] Solve for t\[(\ln (.8)/\ln(.5))*5730=t=1845 yrs\]

OpenStudy (anonymous):

Thank you!